# Variations of metric that preserve a Riemannian submersion and geometry of its fibers

### v1

@@ 4.3 Sectional curvature
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal and $\{V_{n+1},\ldots,V_{n+p}\}$ are Killing fields on $(M,g_{0})$. Then for all $U\in\mathfrak{X}_{\ker\pi_{*}}$ we have
$\partial_{t}\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},U)$
$=\frac{1}{2}\sum\nolimits_{j=n+1}^{n+p}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)\sum\nolimits_{k=n+1}^{n+p}g_{N}([W_{i},W_{j}],W_{k})g_{0}(V_{k},U)$
@@
$=\frac{1}{2}\sum\nolimits_{j,k,l=n+1}^{n+p}g_{N}([W_{i},W_{j}],W_{k})g_{N}([W_{i},W_{j}],W_{l})g_{0}(V_{k},U)g_{0}(V_{l},U)$
$\quad+\sum\nolimits_{j=n+1}^{n+p}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)\sum\nolimits_{k=n+1}^{n+p}g_{N}([W_{i},W_{j}],W_{k})g_{0}(\partial_{t}V_{k},U).$
@@
$4\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},U)=\sum\nolimits_{j=n+1}^{n+p}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)^{2}.$ (52)
From (24), integrability of $\ker\pi_{*}$ and $V_{i}$ being a Killing field, we have
$g_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)=g_{0}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)$
$=g_{0}(\nabla^{0}_{V_{i}}P_{{\cal H}}^{t}\pi^{*}W_{j},U)-g_{0}(\nabla^{0}_{P_{{\cal H}}^{t}\pi^{*}W_{j}}V_{i},U)$
$=-g_{0}(P_{{\cal H}}^{t}\pi^{*}W_{j},\nabla^{0}_{V_{i}}U)-g_{0}(\nabla^{0}_{P_{{\cal H}}^{t}\pi^{*}W_{j}}V_{i},U)$
$-g_{0}(\nabla^{0}_{U}V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j})+g_{0}(\nabla^{0}_{U}V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j})$
$=-(\mathcal{L}_{V_{i}}g_{0})(P_{{\cal H}}^{t}\pi^{*}W_{j},U)+g_{0}([U,V_{i}],P_{{\cal H}}^{t}\pi^{*}W_{j})$
$=-(\mathcal{L}_{V_{i}}g_{0})(P_{{\cal H}}^{t}\pi^{*}W_{j},U)=0.$
Hence, from (4.3) we obtain
$\partial_{t}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)=\sum\nolimits_{k=n+1}^{n+p}g_{N}([W_{i},W_{j}],W_{k})g_{t}(V_{k},U).$ (53)
Using (52) and (53) we obtain (4.12) and (4.12). ∎
Remark 4.13.
We note that for a local $g_{N}$-orthonormal frame $\{W_{n+1},\ldots W_{n+p}\}$ on $N$ we can define the scalar horizontal curvature $K_{{\cal H}(t)}=\sum\nolimits_{i,j=n+1}^{n+p}\sec(P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j})$. By (4.3), we have $K_{{\cal H}}=K_{N}-\frac{3}{4}\|{\tilde{T}}_{t}\|^{2}$, where $K_{N}$ is the scalar curvature of $(N,g_{N})$ and $\|{\tilde{T}}_{t}\|^{2}=\sum\nolimits_{i=n+1}^{n+p}g_{t}(P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],P_{{\cal V}}^{t}[P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}])$. Thus, varying $K_{{\cal H}(t)}$ is equivalent to varying $\|{\tilde{T}}_{t}\|^{2}$. Examples of critical points of functional $g_{t}\mapsto\int_{M}\|{\tilde{T}}_{t}\|^{2}\,\operatorname{vol}_{g_{t}}$ on a domain of a Riemannian submersion $\pi:(M,g)\rightarrow(N,g_{N})$ are metrics, for which $(M,g)$ is a $3$-Sasakian (which was proved in [14]), or a $f$-$K$-contact manifold (which can be proved by similar computations as those in [16]), in each case with vertical characteristic vector fields.

### v2

@@ 5.2 Vertizontal curvature of totally geodesic fibers
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds, where $\{W_{n+1},\ldots,W_{n+p}\}$ are $g_{N}$-orthonormal and $\{V_{n+1},\ldots,V_{n+p}\}$ are bounded vertical vector fields whose restrictions to every fiber ${\cal F}_{x}$ are Killing fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$. Let the fibers of $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be totally geodesic. Then for all $U\in\mathfrak{X}_{{\cal V}}$ we have
$\partial_{t}\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},U)$ $=$ $\frac{1}{2}\sum\nolimits_{j=n+1}^{n+p}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)\cdot$ (57)
$\cdot\big{(}\sum\nolimits_{k=n+1}^{n+p}g_{N}([W_{i},W_{j}],W_{k})g_{0}(V_{k},U)$
$-g_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]+[P_{{\cal H}}^{t}\pi^{*}W_{i},V_{j}],U)\big{)}.$
@@
$=\frac{1}{2}\sum\nolimits_{j=n+1}^{n+p}\big{(}\sum\nolimits_{l=n+1}^{n+p}g_{N}([W_{i},W_{j}],W_{l})g_{0}(V_{l},U)-g_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}]$
$+[P_{{\cal H}}^{t}\pi^{*}W_{i},V_{j}],U)\big{)}^{2}-\sum\nolimits_{j=n+1}^{n+p}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)\cdot\big{(}g_{t}([\partial_{t}V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)$
$+g_{t}([\partial_{t}V_{j},P_{{\cal H}}^{t}\pi^{*}W_{i}],U)\big{)}.$ (58)
@@
$4\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},U)=\sum\nolimits_{j=n+1}^{n+p}g_{t}([P_{{\cal H}}^{t}\pi^{*}W_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)^{2}$ (59)
and from (5.1) we obtain (57). Using (25), (2) and (38) we obtain
$\partial_{t}g_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)=B_{t}([V_{i},P_{{\cal H}}^{t}\pi^{*}W_{j}],U)-g_{t}([V_{i},V_{j}],U)=-g_{t}([V_{i},V_{j}],U).$ (60)
Applying the above together with $V_{i},V_{j}]=-[V_{j},V_{i}]$ in (57) yields (5.5). ∎
Corollary 5.6.
With the assumptions of Theorem 5.5, if $\partial_{t}V_{i}=0$ for all $i\in\{n+1,\ldots,n+p\}$, then $\partial^{2}_{tt}\sec_{M}(P_{{\cal H}}^{t}\pi^{*}W_{i},U)\geq 0$ for all $i\in\{n+1,\ldots,n+p\}$.
Corollary 5.7.
Let $\dim N\geq 2$, let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with totally geodesic fibers and non-integrable horizontal distribution, such that every fiber ${\cal F}_{x}$ is spanned by orthonormal vertical vector fields $\{E_{1},\ldots,E_{n}\}$, whose restrictions to ${\cal F}_{x}$ are Killing vector fields on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$.
For $W\in T_{x}N$, let $\pi^{*}W$ be the $g_{0}$-horizontal lift of $W$. If there exists a fiber ${\cal F}_{y}$ such that for some $X,Y\in T_{y}N$:
  • $1^{\circ}$
    $g_{0}([\pi^{*}X,\pi^{*}Y],E_{1})$ is not constant on ${\cal F}_{y}$, or
  • $2^{\circ}$
    $g_{0}([\pi^{*}X,\pi^{*}Y],E_{1})\neq 0$ on ${\cal F}_{y}$ and there exists a Killing vector field $U_{y}$ on $({\cal F}_{y},g_{0}|_{{\cal F}_{y}})$ such that $g_{0}(U_{y},E_{1})$ is not constant on ${\cal F}_{y}$,
then there exists a metric $g_{s}$ such that $\pi:(M,g_{s})\rightarrow(N,g_{N})$ is a Riemannian submersion with totally geodesic fibers and sectional curvature $\sec_{M}(P_{{\cal H}}^{s}\pi^{*}X,E_{1})$ not constant on the fiber ${\cal F}_{y}$.
Proof.
Let $\cal O$} be an open neighbourhood of $y$ on which there exist $g_{N}$-orthonormal vector fields $\{W_{n+1},\ldots,W_{n+p}\}$ such that at the point $y$ we have: $W_{n+1}=X$, $W_{n+2}=Y$ and $W_{i},W_{j}]=0$ for all $i,j\in\{n+1,\ldots,n+p\}$. Let $f\in C^{\infty}(N)$ be such that $Y(f)\neq 0$ and $W_{j}(f)=0$ at $y$ for all $j\in\{n+1,n+3,\ldots,n+p\}$.
If case $1^{\circ}$ holds, let $\xi=E_{1}$. If case $2^{\circ}$ holds, let $\xi=U$, where $U$ is a vector field whose restriction to a fiber ${\cal F}_{x}$ is a Killing vector field on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ for all $x\in{\cal O}$, and such that $U=U_{y}$ on ${\cal F}_{y}$ ($U_{y}$ can be extended to such vector field $U$ by Remark 3.13).
Let $V_{n+1}=(f\circ\pi)\cdot\xi$ and let $V_{n+2}=\ldots=V_{n+p}=0$ on $\cal O$}. Let $\{{\tilde{g}}_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation such that $\pi:(\pi^{-1}({\cal O}),{\tilde{g}}_{t})\rightarrow({\cal O},g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds on $\pi^{-1}({\cal O})$, with $\{W_{n+1},\ldots,W_{n+p}\}$ and $\{V_{n+1},\ldots,V_{n+p}\}$ as above. Let $\rho\in C^{\infty}(N)$ be a non-zero, compactly supported in $\cal O$} function, such that $\rho=1$ on some neighbourhood of $y$. From now on we will consider the variation $g_{t}$ with $\partial_{t}g_{t}=(\rho\circ\pi)\cdot\partial_{t}{\tilde{g}}_{t}$. Then $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with totally geodesic fibers for all $t\in(-\epsilon,\epsilon)$, by Theorem 3.12.
We note that
$g_{0}([\xi,P_{{\cal H}}^{0}\pi^{*}W_{j}],E_{1})=(\mathcal{L}_{\xi}g_{0})(P_{{\cal H}}^{0}\pi^{*}W_{j},E_{1}),$
since $\xi(g_{0}(P_{{\cal H}}^{0}\pi^{*}W_{j},E_{1}))=0$ by $E_{1}$ being vertical and $g_{0}([\xi,E_{1}],P_{{\cal H}}^{0}\pi^{*}W_{j})=0$ by the vertical distribution being integrable.
From (57) we obtain on ${\cal F}_{y}$:
$\partial_{t}\sec_{M}(P_{{\cal H}}^{t}\pi^{*}X,E_{1})|_{t=0}$
$=-\frac{1}{2}\sum\nolimits_{j=n+2}^{n+p}g_{0}([P_{{\cal H}}^{0}\pi^{*}W_{n+1},P_{{\cal H}}^{0}\pi^{*}W_{j}],E_{1})\cdot g_{0}([V_{n+1},P_{{\cal H}}^{0}\pi^{*}W_{j}],E_{1})$
$=-\frac{1}{2}\sum\nolimits_{j=n+2}^{n+p}g_{0}([\pi^{*}W_{n+1},\pi^{*}W_{j}],E_{1})\cdot g_{0}([(f\circ\pi)\cdot\xi,\pi^{*}W_{j}],E_{1})$
$=-\frac{1}{2}(f\circ\pi)\cdot\sum\nolimits_{j=n+2}^{n+p}g_{0}([\pi^{*}W_{n+1},\pi^{*}W_{j}],E_{1})\cdot g_{0}([\xi,\pi^{*}W_{j}],E_{1})$
$\quad+\frac{1}{2}\sum\nolimits_{j=n+2}^{n+p}g_{0}([\pi^{*}W_{n+1},\pi^{*}W_{j}],E_{1})\cdot g_{0}(\xi,E_{1})\cdot(\pi^{*}W_{j})(f\circ\pi)$
$=-\frac{1}{2}(f\circ\pi)\sum\nolimits_{j=n+2}^{n+p}g_{0}([\pi^{*}W_{n+1},\pi^{*}W_{j}],E_{1})\cdot(\mathcal{L}_{\xi}g_{0})(\pi^{*}W_{j},E_{1})$
$\quad+\frac{1}{2}g_{0}([\pi^{*}X,\pi^{*}Y],E_{1})\cdot g_{0}(\xi,E_{1})\cdot Y(f).$ (61)
With the assumptions about $\xi$, by adjusting $Y(f)$, we can make the right-hand side of (5.2) not constant on ${\cal F}_{y}$. Then there exists $s\in(-\epsilon,\epsilon)$ for which $\sec_{M}(P_{{\cal H}}^{s}\pi^{*}X,E_{1})$ is not constant on ${\cal F}_{y}$. ∎
Example 5.8.
The assumptions of Corollary 5.7 are satisfied e.g., for Hopf fibrations $\pi:(S^{7},g_{0})\rightarrow(S^{4},g_{N})$ and $\pi_{m}:(S^{4m+3},g_{0})\rightarrow(\mathbb{H}P^{m},g_{N})$, $m>1$, with $g_{0}$ being the round metric on the sphere. Then there exist vertical Killing vector fields $\{E_{1},E_{2},E_{3}\}$ on $(M,g_{0})$. We can set $\xi$ in the proof of Corollary 5.7 to be one of the fields $\{E_{1},E_{2},E_{3}\}$, since for all $a\in\{1,2,3\}$, $y\in N$ and $X\in T_{y}N$ there exists $Y\in T_{y}N$ such that the function $g_{0}([\pi^{*}X,\pi^{*}Y],E_{a})$ is not constant on ${\cal F}_{y}$ [9]. We note that on every fiber ${\cal F}_{x}\equiv S^{3}$ there exists also another family of Killing fields $\{\xi_{1},\xi_{2},\xi_{3}\}$ such that functions $g(E_{a},\xi_{b})$ are not constant on ${\cal F}_{x}$ (since the linear space of Killing vector fields on $S^{3}$ has dimension $6$, see e.g., [4, Section 2] for an explicit presentation of those Killing fields and relations between them). In fact, vector fields $\xi_{a}$ on ${\cal F}_{x}$ can be obtained as $P_{{\cal V}}^{0}[\pi^{*}X,\pi^{*}Y]$ for some $X,Y\in T_{x}N$ [10].
Corollary 5.7 indicates existence of fat bundles with vertizontal curvatures not constant along fibers, answering negatively Problem 1 in [20].
We have also the following result about the stability of vertical Killing vector fields with respect to variations preserving a Riemannian submersion and metric induced on its fibers.
Theorem 5.9.
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$ and (12) holds. Let $K$ be a vertical Killing vector field on $(M,g_{0})$, then $K$ is a Killing vector field for all $\{g_{t},t\in(-\epsilon,\epsilon)\}$ if and only if for all $\{V_{n+1},\ldots,V_{n+p}\}$ in (12) we have $V_{i},K]=0$.
Proof.
Since $K$ is a Killing vector field, $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ and $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$, we have for all $t\in(-\epsilon,\epsilon)$, all vertical fields $U_{1},U_{2}$ and all $g_{t}$-horizontal fields $Z_{1},Z_{2}$:
$0=(\mathcal{L}_{K}g_{t})(U_{1},U_{2})=(\mathcal{L}_{K}g_{t})(Z_{1},Z_{2}).$
Let $i\in\{n+1,\ldots,n+p\}$ and let $U$ be a vertical vector field. Then, using (12), integrability of $\mathcal{V}$, (2) and (38) we obtain
$\partial_{t}((\mathcal{L}_{K}g_{t})(P_{{\cal H}}^{t}\pi^{*}W_{i},U))=\partial_{t}\big{(}K(g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i},U))-g_{t}([K,P_{{\cal H}}^{t}\pi^{*}W_{i}],U)$
$-g_{t}([K,U],P_{{\cal H}}^{t}\pi^{*}W_{i})\big{)}$
$=-B_{t}([K,P_{{\cal H}}^{t}\pi^{*}W_{i}],U)-g_{t}([K,\partial_{t}P_{{\cal H}}^{t}\pi^{*}W_{i}],U)$
$=g_{t}([K,V_{i}],U).$
In Example 5.8, if we take $\xi\in\{\xi_{1},\xi_{2},\xi_{3}\}$ then $E_{1},E_{2},E_{3}$ will remain Killing fields on $(S^{4m+3},g_{t})$, since $\xi_{a},E_{b}]=0$ for all $a,b\in\{1,2,3\}$. On the other hand, for all $a\in\{1,2,3\}$ and all $x\in N$ the restriction of $\xi_{a}$ to a fiber ${\cal F}_{x}$ will be a Killing field on $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$, but in general $\xi_{a}$ may not be a Killing field on $(S^{4m+3},g_{t})$ for $t\neq 0$.
The metric properties of $(M,g_{s})$ obtained in Corollary 5.7 may be significantly different than those of $(M,g_{0})$. In particular, metrics obtained in Example 5.8 will not be 33-Sasakian. However, we note that manifolds in Example 5.8 are 33-sphere bundles, which over $S^{4}$ are classified up to isomorphism [7].
Acknowledgements
The author is grateful to the late Professor Paweł Walczak for helpful remarks about an early version of this paper.